OpenStudy (anonymous):

what is the value of k,-4 is a zero of the polynomial x2(x square)-x-(2k+2)?

OpenStudy (anonymous):

i need the procedure

OpenStudy (lgbasallote):

if -4 is a zero that means if you substitute -4 for x..you should get 0 \[x^2 - x - (2k + 2) = 0\] \[\Rightarrow (-4)^2 - (-4) - (2k+2) = 0\] \[\Rightarrow 16 + 4 -(2k + 2) = 0\] \[\Rightarrow 20 - 2k - 2 = 0\] \[\Rightarrow 18 - 2k = 0\] \[\Rightarrow -2k = -18\] i suppose you can do the rest??

OpenStudy (anonymous):

if the zero of d polynomial is given it basically implies dat after poppin dat value in d place of x in2 d the polynomial u basically get 0 as ur result.so (-4)^2-(-4)-(2k+2)=0 so 16+4-2k-2=0,k=9